OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 4x − ___

OpenStudy (anonymous):

write you x-intercepts in factored form and expand it.

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

x intercepts at x = -6 and x = 2, right? so factored form is y= (x-r)(x-s), where r= -6 and s= 2. plug it in then expand it..

OpenStudy (anonymous):

y=(x--6)(x-2)

OpenStudy (anonymous):

so it will be like this: y=(x+6)(x-2).. expand this

OpenStudy (anonymous):

y=x^2-2x+6x-12

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

x^2+4x-12

OpenStudy (anonymous):

there you go, you already have the answer.. just look at it and find the missing term:)

OpenStudy (anonymous):

Can you explain to me what the answer is because im still not so sure @Data_LG2

OpenStudy (anonymous):

y = x2 + 4x − ___.. compare this to your answer above.. the missing term will be ...

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

kind of.. read your question again and it says: (Do not include the negative sign in your answer.) so the answer is?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

a medal for you